### v2

@@ Abstract
Neutrinos are elementary particles that interact only very weakly with matter. Neutrino experiments are therefore usually big, with masses on the multi-ton scale. The thresholdless interaction of coherent elastic scattering of neutrinos on atomic nuclei leads to drastically enhanced interaction rates, that allows for much smaller detectors. The study of this process gives insights into physics beyond the Standard Model of particle physics. The Conus+ experiment was designed to first detect elastic neutrino-nucleus scattering in the fully coherent regime with low-energy neutrinos produced in nuclear reactors. For this purpose, semiconductor detectors based on high-purity germanium crystals with extremely low energy threshold of 160180 eV were developed. Here we show the first observation of a neutrino signal with a statistical significance of 3.7 sigma from the Conus+ experiment, operated at the nuclear power plant in Leibstadt, Switzerland. In 119 days of reactor operation (395106) neutrinos were measured compared to a predicted number from calculations assuming standard model physics of (34759) events. The good agreement between data and prediction constrains many parameters in various theoretical models. With increased precision, there is potential for fundamental discoveries in the future. The Conus+ results in combination with other measurements of this interaction channel might therefore mark a starting point for a new era in neutrino physics.
@@ Introduction
Neutrinos are known for their very tiny interaction rate with matter. This is why they are usually very hard to detect, despite their high abundance and the existence of very strong neutrino sources. The most common detection channels such as the inverse beta decay reaction or neutrino-electron scattering usually require target masses in the ton to kiloton scale. In the Standard Model (SM) of particle physics, neutrinos can couple to quarks by the exchange of a mediating Z boson. For small momentum exchanges, the possibility of coherent scattering of neutrinos on the sum of all nucleons of an atomic nucleus was predicted in 1974 [1]. For this interaction the reaction rate (cross-section) is enhanced by few orders of magnitude as it scales approximately with the squared number of neutrons in the target nucleus. Therefore, it is in principle possible to construct neutrino detectors on the kilogram scale using this channel.
@@
Neutrino sources suitable for the measurement of the coherent elastic neutrino-nucleus scattering (CENS) need to have energies of a few up to tens of MeV. Intense neutrino sources fulfilling this condition are, for example, the Sun, supernova explosions, accelerator-based sources, or nuclear reactors. It took 43 years after its prediction until CENS was first detected in 2017 by the COHERENT experiment in a scintillating crystal of cesium iodide [2]. Here, the neutrinos are generated at a Spallation Neutron Source (SNS) when pions decay at rest. At the same neutrino source the CENS measurement was confirmed with argon [3] and germanium [4] as target materials. Nuclear reactors emit neutrinos of lower energies than SNS sources. This has the advantage of a strongly enhanced sensitivity for several parameters in beyond the standard model (BSM) theories. However, it makes the detection of neutrinos much harder because of the requirement of extremely low energy thresholds in the detector.
@@
There is a long history of successful experiments using reactor antineutrinos as sources, including the first neutrino detection in 1956 [5]. An intense antineutrino flux is produced after the fission process during the beta decay of the fission fragments. Nuclear reactors can be considered as well-defined, pure and point-like sources of specific neutrino flavor type, the electron antineutrinos. In recent years, reactor experiments allowed to study neutrino oscillation parameters [6, 7, 8, 9] and constrain the existence of sterile neutrinos [10, 11, 12, 13]. Today, there is a worldwide effort to measure CENS close to nuclear reactors [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. In terms of neutrino oscillation studies, it is a complementary approach, since CENS is sensitive to all known neutrino flavors, the electron, muon and tau neutrinos/antineutrinos, whereas the inverse beta decay as standard detection technique for reactor neutrinos is only sensitive to electron antineutrinos. In addition, there is a wide range of studies and topics that can be addressed in CENS measurements. For example, they can be used to probe nuclear structures, they are highly relevant in astrophysics, and they also provide important input for present and future dark matter experiments, which are limited in sensitivity by the CENS rate of solar neutrinos. Very recently, first indications for such a CENS signal were found in dark matter experiments with xenon as target [24, 25]. Moreover, Conus+ technology has the potential to use neutrinos for reactor monitoring and safeguard applications in the future.
@@
The Conus experiment [18] began operating in 2018 at the nuclear power plant in Brokdorf, Germany (KBR). There, the experiment took data until 2022 at a distance of only 17.1 m from the center of the reactor core. At the maximum thermal power of the reactor of 3.9 GW, the corresponding flux at the detector position was estimated to be 1013 antineutrinos per cm2 and second. The experimental setup consisted of four high-purity germanium (HPGe) detectors. Each diode had a mass of about 1 kg leading to a total fiducial germanium mass of kg [26]. As a final result of the measurement at KBR, the neutrino flux was constrained with 90% confidence level to a factor 1.6 above the signal expectation [27]. This world-best upper limit on the CENS interaction rate at nuclear reactors so far allowed us to exclude significant deviations from the SM or to test the standard description of signal quenching due to dissipation effects in the germanium material in the energy region of interest.
@@
In 2023, the Conus setup moved to another power plant in Leibstadt, Switzerland (KKL), since the reactor at KBR stopped operation. At KKL the experiment continued as Conus+. Here, antineutrinos are created in a boiling water reactor with a thermal power of 3.6 GW. The setup is placed at a distance of about 20.7 m from the center of the KKL reactor core. Before installation, the HPGe detectors called CC5 were refurbished to improve the energy threshold and the detection efficiency at low energy. In this way, the sensitivity was improved despite a higher level of environmental radioactivity and a slightly lower nominal neutrino flux of 1013 antineutrinos/(cm2s) at the new site. The predicted number of neutrino interactions in the four detectors increased almost an order of magnitude, mainly due to the improved energy threshold and trigger efficiency [28].
@@
The search for CENS in nuclear reactors is a challenging task for various reasons. Since a high neutrino flux is required, positions close to the reactor core inside the reactor building are preferred. The environment inside this inner control zone is quite different from the working conditions in a common research laboratory. There are several restrictions related to the materials allowed, earthquake safety, access, data transfer, etc. Appropriate solutions on all these topics were found in close cooperation with the KBR and KKL staff. In addition, there is limited protection against cosmic radiation, as the overburden at the CONUS+ site corresponds only to the equivalent 7.4 m water. In general, the radioactivity level has to be kept under control to perform successful rare event searches. For example, cosmic muons produce electromagnetic cascades and neutrons in the building structure and shield materials. These cascades can create event signatures similar to neutrinos in the HPGe detectors. Mitigation of detector signals created by such cosmic radiation or environmental radioactivity (background events) is achieved by using an effective shield structure around the detectors [28] as depicted in Fig.1. Furthermore, the energy of the nuclear recoils after neutrino scattering is very low. The unit used for the energy measured by the Ge detectors is given in eV ( J) and should be interpreted as ionization energy. It was a longstanding effort to reach the required threshold levels in the detectors. Our currently lowest threshold level of 160 eV is only two orders of magnitude above the typical semiconductor band gap, which defines the minimum energy to create one electron-hole pair.
@@
The basic concept of antineutrino detection in Conus+ is to measure an energy spectrum under stable conditions in phases for which the reactor is running or stopped (reactor on and off). Most events originate from cosmic radiation, which is fully independent of the reactor operation condition. During on-phases additional neutrino events on top of the cosmogenic background contribution are expected with a characteristic spectral shape. The neutrino signal can be extracted from a comparison of the data sets with the reactor on and off. The reactor off-phases without neutrino signal are obtained in maintenance periods for refueling, which are typically once per year with a duration of approximately 1 month.
The energy spectrum measured between [] keV during reactor on period is depicted in Fig.2, together with the expected background calculated using a well-validated GEANT4 framework [29]. A good understanding of the background composition is mandatory for our neutrino analysis. Following the experience gained at the Brokdorf reactor, the background spectra were adjusted to the new site in Leibstadt. The contributions of muon-induced, neutron and gamma components were validated in a dedicated measurement campaign with multiple detector technologies before installing the Conus+ setup [30]. Uncorrelated background contributions, which are independent of the reactor thermal power condition, are measured in the reactor off-phase.
The dominant background contributions, such as the muon-induced or neutron background, are identical for the three detectors used in the analysis. A comparison between background model and data is shown in Fig.2. In addition to the neutrons, which are created by muons in the materials of the Conus+ shield, there is a relevant fraction of direct neutrons up to 100 MeV energy already produced in the atmosphere in cosmic ray air showers. The background during reactor off was lower since a vessel lid with a thickness of a few cm of steel was placed right above the Conus+ room during reactor outage providing additional overburden. According to simulations, this reduction was 3% for muons and 19% for direct cosmic neutrons. Other components contributing to the background spectrum are from metastable states in germanium and the radioactive noble gas radon. Variations in the radon level inside the detector chamber were corrected. Otherwise, no significant differences are observed during reactor on and off. The reactor correlated background events, in particular the ones from neutrons produced in the reactor, are thoroughly studied in [30] and found to be negligible.
@@
The data set used in the analysis reported here includes reactor on periods between November 2023 and July 2024 (327 kg d) and a off period during reactor outage in May 2024 (60 kg d). The data sets of the three detectors are fitted simultaneously in an energy window between 160 eV and 800 eV. The signal is extracted based on a profile likelihood ratio test. Systematic uncertainties are considered with Gaussian pull terms. The data acquisition system in principle allows for background rejection by studying the shape of digitized pulses [35]. This option was not yet applied in the current analysis, but it is planned to be used in future analyses.
From the combined fit, a neutrino signal of () events in the reactor on data set was obtained. The significance of this event excess corresponds to . The neutrino signal at low energies of the spectrum is illustrated in Fig.3. Fits using just single detectors independently give consistent results. The result was cross-checked in two implementations of the likelihood fit and agreed within less than 2%. Additional systematic uncertainties related to background model, non-linearity correction and fit systematics were studied independently and are included in the fit result.
@@
A high-statistics CENS measurement might open a new phase in fundamental physics and will allow to study physics within and beyond the SM. The measured CENS rate can for example be affected by new mediator particles similar to the Z boson, electromagnetic properties of neutrinos, or non-standard interactions. Moreover, it is possible to study the Weinberg angle at low energies, the existence of sterile neutrinos or supernovae astrophysics. There are also connections to dark matter experiments. A precise CENS measurement will allow to learn more about the neutrino sources as the Sun or nuclear reactors. The evolution of the reactor thermal power and fissile isotope concentrations in fuel elements could be monitored with rather small and mobile neutrino detectors. In summary, there is a wide range of topics ranging from BSM theories, nuclear physics and astrophysics that can be addressed with CENS measurements at nuclear reactors. The Conus and Conus+ experiments are pioneering in this field.
@@ Acknowledgments
We thank all divisions and workshops involved at the Max-Planck-Institut für Kernphysik in Heidelberg to set up the CONUS+ experiment, in particular T. Apfel, M. Reissfelder, T. Frydlewicz and J. Schreiner. We also thank Mirion Technologies (Canberra) in Lingolsheim for the detector upgrades and their highly professional support. Our deepest gratitude goes to the Leibstadt AG for hosting and supporting the CONUS+ experiment with special thanks to P. Graf, P. Kaiser, L. Baumann, R. Meili and A. Ritter.
@@
The evolution of the main environmental parameters during reactor on and off is shown in figure 4-left, figure 4-center and figure 4-right, for the C5, C2, and C3 detectors, respectively. The excluded data periods are depicted (red-shadow area). The shape of the noise peak reconstructed in the lowest channels of the data acquisition system (DAQ) was found to follow a Gaussian distribution [28] and its width (FWHM) was monitored over time with variations below 2 eVee (electron equivalent energy). The data set was selected to ensure that the noise rate variations are below 20%. Variations in the cryocooler power are not solely caused by changes in room temperature. They were also induced by instabilities in the recirculation of the cooling liquid, an issue that was solved for run 2 in Conus+.
@@
The trigger efficiency was determined by injecting artificial signals produced by a pulse generator with the same rise time as the physical signals. A detailed scan allowed us to measure the detector response as a function of the energy. The description of the experimental trigger efficiency curves is depicted in [28] and was obtained using the fit function:
with being a value in [99, 109] and in [51, 54], respectively. In particular, the trigger efficiency remains over 90% down to 140 eVee for all detectors. The evolution of the trigger efficiency curve parameters during run 1 was studied with different measurements, remaining stable with differences of less than 2% throughout the run.
@@
Proper energy calibration is fundamental for achieving a precise neutrino measurement. The energy was calibrated using the binding energies of the K shell (10.37 keV) and L shell (1.30 keV) from the decays of 71Ge inside the HPGe diodes, considering a linear behavior in this energy range. The 71 Ge is created by the capture of neutrons up to a few MeV energy. The half-life of 71Ge is 11.4 d, so a few weeks of data collection is enough to record most of the decays of 71Ge after activation. The 71Ge radioisotope is also continuously produced at shallow depth by neutrons released in muon interactions with Pb of the Conus+ shield [28]. Thus, it is possible to monitor the stability of the energy scale during the whole measurement of this in situ activation, observing variations below 2%. A specific 252Cf irradiation was performed at the end of run 1, collecting in 45 d of the measurement more than 5000 (700) events from the K (L) shell in each detector. In this way, an energy calibration uncertainty below 5 eVee is achieved. The HPGe diode also produces 158 eV X-rays from the binding energy of the M shell of the 71Ge decays. Using the ratio of the K and L shells, a total of 100 events per detector was expected after irradiation with 252 Cf. Although an indication of such events is seen, no conclusive signal was yet observed due to the proximity to the noise edge and the lack of statistics. Future irradiation after upgrades is planned to measure the M shell line in the Conus+ detectors.
@@
The linearity of the DAQ chain in the sub-keVee region was investigated by injecting artificial pulser signals. The results are shown in figure 5. Deviations from a pure linear behavior are observed below 250 eVee, where a few eVee variations have a strong impact on the CENS signal. They were attributed to two nearly independent DAQ-related effects. First, as the amplitude of the signals decreases, artificial delays of the trigger time stamp were observed up to a few s. Due to the relatively short shaping time of 4 s, these delays affect the reconstructed energy value obtained by trapezoidal shaping. The corresponding negative non-linearity effect reaches its maximum of approximately 15 eVee at 120 eVee. Second, the drop in trigger efficiency below 120 eVee is responsible for an artificial positive deviation from linearity, overwhelming the above-mentioned effect. This effect was also observed in [34]. The energy non-linearity was corrected in the CENS analysis and was taken into account during the event energy reconstruction.
@@
The energy resolution of the detectors at low energies was evaluated using artificial signals generated by a pulse generator with the same pulse rise time as the physical signals. The pulses are injected through a specific circuit implemented in the HPGe preamplifier. For the C2 detector, a pulser resolution of () eVee (FWHM) was found, while the C5 and C3 detectors have a pulser resolution of () eVee (FWHM). An additional contribution to the total energy resolution comes from the statistical fluctuation of the number of electron-hole pairs produced in the Germanium crystals in the case of an event. This contribution is
@@
with = 2.96 eV, the average energy needed to create a single electron-hole pair in Germanium, and the fano factor F. For this analysis a value of F = 0.1096 was used.
@@
The spectrum in the HPGe detectors is dominated by background events (mainly produced by muon-induced neutrons). Selection cuts are applied to reduce the background while keeping the CENS signal. Four different selection cuts are applied to the data: First, the muon-veto system allows for efficient suppression of the impact of cosmic radiation, using a 450 s anti-coincidence window between the veto and HPGe signals. The average rate detected in the muon veto during reactor ON is () Hz and decreases to () Hz in reactor off periods. For this reason, the dead time induced by the muon veto system was independently estimated for both periods with average values of 12.3% (reactor on) and 9.6% (reactor off). This cut successfully suppressed the background with a rejection efficiency of more than 99% in most of the energy range.
Second, inhibit signals are generated when the increasing baseline has reached saturation of the dynamic range of the transistor reset preamplifier (TRP) for each HPGe detector [28]. An anti-coincidence window of 1 to 2.5 ms (depending on the detector) is applied to veto unwanted spurious HPGe detector signals generated shortly after the resets. This cut suppressed 30% events after the muon veto anti-coincidence at low energy, becoming negligible above 5 keVee. The dead time induced by this cut is calculated combined with the previously mentioned muon-veto dead time to avoid the overlapping of both veto windows. An additional dead time of 0.5-2.1% (depending on the detector) is estimated.
Third, the time difference distribution (TDD) of events is studied in each channel as proposed in [26]. In this way, noise events from failures in the pile-up cut of the DAQ and from microphonics are efficiently removed with nearly no loss in live time.
Finally, an anti-coincidence cut is applied between different HPGe detectors with a 5 ms time window. The probability of a neutrino interacting with different detectors is negligible, while for other backgrounds, such as muon-induced neutrons created in the shield, simultaneous hits in several detectors at once can be expected. This cut suppresses about 5% of the background after the muon veto anti-coincidence in the [0.4-1] keVee energy range without inducing a significant amount of dead time. The measured suppression efficiency is in good agreement with the background simulations.
[0.4-1.0] keVee | ||||||
Detector | C5 | C2 | C3 | |||
Rector period | On | Off | On | Off | On | Off |
muon veto anti. | 99.3% | 99.3% | 99.8% | 99.8% | 99.8% | 99.8% |
TRP anti. | 35.9% | 39.6% | 43.4% | 44.0% | 43.9% | 46.5% |
TDD | 0.1% | 0.1% | 0.2% | 0.5% | 0.2% | 0.1% |
neutron anti. | 6.8% | 6.6% | 4.3% | 4.8% | 7.2% | 6.8% |
combined | 99.6% | 99.6% | 99.9% | 99.9% | 99.9% | 99.9% |
[2-8] keVee | ||||||
Detector | C5 | C2 | C3 | |||
Rector period | On | Off | On | Off | On | Off |
muon veto anti. | 98.9% | 98.9% | 98.6% | 98.6% | 98.7% | 98.9% |
TRP anti. | 19.7% | 25.0% | 3.9% | 5.5% | 6.1% | 6.7% |
TDD | 0.1% | 0.1% | 0.1% | 0.1% | 0.1% | 0.1% |
neutron anti. | 6.5% | 7.0% | 4.3% | 4.3% | 6.7% | 6.5% |
combined | 99.2% | 99.2% | 99.1% | 99.1% | 98.9% | 99.1% |
[…]
The rejection efficiencies of these selection cuts during reactor on and off are summarized in table 2 for the C5, C2, and C3 detectors in different energy regions. A total rejection efficiency over 99% is achieved in the [0.4-1] keVee energy range for all detectors. The total dead time induced by the selection cuts is dominated by the muon veto anti-coincidence with values between 12.8% and 14.4% in the reactor on periods. In addition, the average dead time induced by the DAQ is calculated to be below 2%.
@@
The background model used in the analysis of the Conus+ data is based on Monte Carlo simulations using the Geant4-based framework MaGe [39]. With this framework, three categories of possible backgrounds were investigated following the approach taken in [29]:
1.
Cosmogenic components (Muons and neutrons)
2.
Natural radioactivity (external and internal to shield)
3.
Artificial components (Reactor-correlated components, surface contaminations)
In general, a full decomposition of the background in both, reactor on and off data was achieved for all three detectors used in the analysis. In the following section, the major sources of background will be described in more detail.
@@
As described in [30], the impact of cosmogenic neutrons with energies up to 100 MeV was investigated by first propagating the expected neutron spectrum at the KKL location (based on [40],[41]) with a flux of neutrons s-1cm-2 through a model of the reactor building. Then the resulting flux is tracked inside the Conus+ room. This flux was used as the basis for a next simulation, in which neutrons were started isotropically from a half-sphere around the Conus+ shield and their contribution to the Conus+ background was measured. The simulations show a contribution of counts d-1kg-1 in each detector in the energy range between 0.4 keVee and 1 keVee, which corresponds to approximately half of the background counts in this region, respectively.

### v3

@@ Abstract
Neutrinos are elementary particles that interact only very weakly with matter. Neutrino experiments are, therefore, usually big, with masses in the multi-tonne range. The thresholdless interaction of coherent elastic scattering of neutrinos on atomic nuclei leads to greatly enhanced interaction rates, which allows for much smaller detectors. The study of this process gives insights into physics beyond the Standard Model of particle physics. The CONUS+ experiment was designed to first detect elastic neutrino–nucleus scattering in the fully coherent regime with low-energy neutrinos produced in nuclear reactors. For this purpose, semiconductor detectors based on high-purity germanium crystals with extremely low-energy thresholds were developed. Here we report the first observation of a neutrino signal with a statistical significance of 3.7 from the CONUS+ experiment, operated at the nuclear power plant in Leibstadt, Switzerland. In 119 days of reactor operation () neutrinos were measured compared with a predicted number from calculations assuming Standard Model physics of () events. With increased precision, there is potential for fundamental discoveries in the future. The CONUS+ results in combination with other measurements of this interaction channel might therefore mark a starting point for a new era in neutrino physics.
@@ Introduction
Neutrinos are known for their tiny interaction rate with matter. This is why they are usually very hard to detect, despite their high abundance. The most common detection channels such as the inverse beta decay reaction or neutrino-electron scattering usually require target masses in the ton to kiloton scale. In the Standard Model (SM) of particle physics, neutrinos can couple to quarks by the exchange of a mediating Z boson. For small momentum exchanges, the possibility of coherent scattering of neutrinos on the sum of all nucleons of an atomic nucleus was predicted in 1974 [1]. For this interaction the reaction rate (cross-section) is enhanced by few orders of magnitude as it scales approximately with the squared number of neutrons in the target nucleus. Therefore, it is in principle possible to construct neutrino detectors on the kilogram scale using this channel.
@@
It took 43 years after its prediction until CENS was first detected in 2017 by the COHERENT experiment in a scintillating crystal of cesium iodide [2]. Later, COHERENT confirmed the measurement with argon [3] and germanium [4] as target materials. Here, the neutrinos are generated at a Spallation Neutron Source (SNS) when pions decay at rest. A complementary approach for CENS detection is to use nuclear reactors as a source. An advantage is the lower neutrino energies as compared to the SNS source, which offers an enhanced sensitivity for several parameters in beyond the standard model (BSM) theories. At higher neutrino energies above approximately 10 MeV, there is a transition from the fully to the partially coherent regime, and uncertainties related to the nuclear structure become relevant. Measurements at reactor can be used to get a clean coherent scattering signal and set an anchor point for the cross-section. With this knowledge, experiments such as COHERENT can study the nuclear structure of the corresponding target isotope with higher-energy neutrinos. Nuclear reactors provide a very high flux of pure electron antineutrino, whereas the accelerator-based SNS source involves several types of neutrinos, including muon neutrinos and antineutrinos.
@@
There is a long history of successful experiments using reactor antineutrinos as sources, including the first neutrino detection in 1956 [5]. In recent years, reactor experiments allowed to study neutrino oscillation parameters [6, 7, 8, 9] and constrain the existence of sterile neutrinos [10, 11, 12, 13]. Today, there is a worldwide effort to measure CENS near nuclear reactors [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. In terms of neutrino oscillation studies, it is a complementary approach, since the CENS process is sensitive to the three known neutrino flavors, whereas the inverse beta decay as a standard detection technique is only sensitive to electron antineutrinos. There is a wide range of studies that can be addressed in CENS measurements. For example, they are highly relevant for probing nuclear structures, for astrophysical studies, and for present and future dark matter experiments, which are limited in sensitivity by the CENS rate of solar neutrinos. Recently, indications for a CENS signal were found in dark matter experiments with xenon as target [24, 25]. Moreover, Conus+ technology has the potential to use neutrinos for reactor monitoring and safeguard applications in the future.
@@
The Conus experiment [18] began operating in 2018 at the nuclear power plant in Brokdorf (KBR), Germany, and took data until 2022. The experimental setup consisted of four high-purity germanium (HPGe) detectors. Each diode had a mass of about 1 kg leading to a total fiducial germanium mass of kg [26]. As a final result of the measurement at KBR, the neutrino flux was constrained with 90% confidence level to a factor 1.6 above the signal expectation [27]. This world-best upper limit on the CENS interaction rate at nuclear reactors so far allowed us to exclude significant deviations from the SM or to test the standard description of signal quenching due to dissipation effects in the germanium material in the energy region of interest.
@@
In 2023, the Conus setup moved to another power plant in Leibstadt, Switzerland (KKL), since the reactor at KBR stopped operation. At KKL the experiment continued as Conus+. Here, antineutrinos are created in a boiling water reactor with a thermal power of 3.6 GW. The setup is placed at a distance of about 20.7 m from the center of the KKL reactor core. Before installation, the HPGe detectors called CC5 were refurbished to improve the energy threshold and the detection efficiency at low energy. In this way, the sensitivity was improved despite a higher level of environmental radioactivity and a slightly lower nominal neutrino flux of 1013 antineutrinos/(cm2s) at the new site. The predicted rate of neutrino interactions increased almost an order of magnitude, mainly due to the improved energy threshold and trigger efficiency [28].
@@
The search for CENS in nuclear reactors is a challenging task for various reasons. Since a high neutrino flux is required, positions close to the reactor core inside the reactor building are preferred. The environment inside this inner control zone is quite different from the working conditions in a common research laboratory. There are several constraints related to material restrictions, earthquake safety, access and data transfer. Appropriate solutions on all these topics were found in close cooperation with the KBR and KKL staff. In addition, there is limited protection against cosmic radiation, as the overburden at the CONUS+ site corresponds only to the equivalent 7.4 m water. In general, the radioactivity level has to be kept under control to perform successful rare event searches. For example, cosmic muons produce electromagnetic cascades and neutrons in the building structure and shield materials. These cascades can create event signatures similar to neutrinos in the HPGe detectors. Mitigation of detector signals created by such cosmic radiation or environmental radioactivity (background events) is achieved by using an effective shield structure around the detectors [28] as depicted in Fig.1. Furthermore, the energy of the nuclear recoils after neutrino scattering is very low. The unit used for the energy measured by the Ge detectors is given in eV ( J) and should be interpreted as ionization energy. It was a longstanding effort to reach the required threshold levels in the detectors. Our currently lowest threshold level of 160 eV is only two orders of magnitude above the typical semiconductor band gap, which defines the minimum energy to create one electron-hole pair.
@@
The basic concept of antineutrino detection in Conus+ is to measure an energy spectrum under stable conditions in phases for which the reactor is running or stopped (reactor on and off). Most events originate from cosmic radiation, which is fully independent of the reactor operation condition. During on-phases additional neutrino events are expected with a characteristic spectral shape. The reactor off-phases without neutrino signal are obtained in maintenance periods for refueling, which are typically once per year with a duration of approximately 1 month.
The energy spectrum during reactor on period is depicted in Fig.2, together with the expected background calculated using a well-validated GEANT4 framework [29]. A good understanding of the background composition is mandatory for our neutrino analysis. Following the experience gained at KBR, the background spectra were adjusted to the new KKL site. The contributions of muon-induced, neutron and gamma components were validated in a dedicated measurement campaign with multiple detector technologies before installing the Conus+ setup [30]. Uncorrelated background contributions, which are independent of the reactor thermal power condition, are measured in the reactor off-phase.
a
b
The dominant background contributions, such as the muon-induced or neutron background, are identical for the three detectors used in the analysis. In addition to the neutrons, which are created by muons in the materials of the Conus+ shield, there is a relevant fraction of direct neutrons up to 100 MeV energy already produced in the atmosphere in cosmic ray air showers. The background during reactor off was lower since a vessel lid with a thickness of a few cm of steel was placed right above the Conus+ room during reactor outage providing additional overburden. According to simulations, this reduction was 3% for muons and 19% for direct cosmic neutrons. Other components contributing to the background spectrum are from metastable states in germanium and the radioactive noble gas radon. Variations in the radon level inside the detector chamber were corrected. Otherwise, no significant differences are observed during reactor on and off. The reactor correlated background events, in particular the ones from neutrons produced in the reactor, are thoroughly studied in [30] and found to be negligible.
@@
The data set used in the analysis reported here includes reactor on periods between November 2023 and July 2024 (327 kg d) and an off period during reactor outage in May 2024 (60 kg d). The data sets of the three detectors are fitted simultaneously in an energy window between 160 eV and 800 eV. The signal is extracted based on a profile likelihood ratio test. Systematic uncertainties are considered with Gaussian pull terms. The data acquisition system in principle allows for background rejection by studying the shape of digitized pulses [35]. This option was not yet applied in the current analysis, but it is planned to be used in future analyses.
From the combined fit, a neutrino signal of () events in the reactor on data set was obtained. The significance of this event excess corresponds to . The neutrino signal at low energies of the spectrum is illustrated in Fig.3. Fits using just single detectors independently give consistent results. The result was cross-checked in two implementations of the likelihood fit. Additional systematic uncertainties related to background model, non-linearity correction and fit systematics were studied independently and are included in the final result.
a
b
c
@@
A high-statistics CENS measurement might open a new phase in fundamental physics and will allow to study physics within and beyond the SM [38]. The measured CENS rate can, for example, be affected by new mediator particles similar to the Z boson, electromagnetic properties of neutrinos, or non-standard interactions. Moreover, it is possible to study the Weinberg angle at low energies, the existence of sterile neutrinos, or supernovae astrophysics. With a precise CENS measurement, more can be learned about neutrino sources as the Sun or nuclear reactors. The evolution of reactor thermal power and fissile isotope concentrations in fuel elements could be monitored with rather small and mobile neutrino detectors. In summary, there is a wide range of topics ranging from BSM theories, nuclear physics, and astrophysics that can be addressed with CENS measurements at nuclear reactors. The Conus and Conus+ experiments are pioneering in this field.
@@ Acknowledgments
We thank all divisions and workshops involved at the Max-Planck-Institut für Kernphysik in Heidelberg to set up the CONUS+ experiment, in particular T. Apfel, M. Reissfelder, T. Frydlewicz and J. Schreiner. The authors also thank Mirion Technologies (Canberra) in Lingolsheim for the detector upgrades and their highly professional support. Our deepest gratitude goes to Kernkraftwerk Leibstadt AG for hosting and supporting the CONUS+ experiment with special thanks to P. Graf, P. Kaiser, L. Baumann, R. Meili and A. Ritter. All authors are members of the CONUS Collaboration.
@@
The evolution of the main environmental parameters during reactor on and off is shown in Extended Data Fig. 4 for the three detectors. The shape of the noise peak reconstructed in the lowest channels of the data acquisition system (DAQ) was found to follow a Gaussian distribution [28] and its width (FWHM) was monitored over time with variations below 2 eVee (electron equivalent energy). The data set was selected to ensure that the noise rate variations are below 20%.
In a dedicated study, we confirmed the correlation between noise rate and cryocooler power. Stability was improved compared to the CONUS setup in Brokdorf (KBR) by replacing the 2-fan ventilation system with a water-cooled chiller system for the pulse tube cryocoolers [28]. Even at cryocooler power variations of up to 30 W no correlation was observed with the count rate in the region of interest above 160 eVee.
@@
The trigger efficiency was determined by injecting artificial signals produced by a pulse generator with the same rise time as the physical signals. The pulses are injected through a specific circuit implemented in the HPGe preamplifier. A detailed scan allowed us to measure the detector response as a function of the energy. [28] The trigger efficiency remains over 90% down to 140 eVee for all detectors. The evolution of the trigger efficiency curve parameters during run 1 was studied with different measurements, remaining stable with differences of less than 2% throughout the run.
@@
The energy was calibrated using the binding energies of the K-shell (10.37 keV) and L-shell (1.30 keV) from the decays of 68Ge/71Ge inside the HPGe diodes, considering a linear behavior in this energy range. At shallow depth, the 68Ge/71Ge radioisotopes are continuously produced by cosmic and muon-induced neutrons [28]. Thus, it is possible to monitor the stability of the energy scale during the whole measurement of this in situ activation, observing variations below 2%. A specific 252Cf irradiation was performed at the end of run 1, collecting in 45 d of the measurement more than 5000 (700) events from the K (L) shell in each detector. In this way, an energy calibration uncertainty below 5 eVee is achieved. The HPGe diode also produces 158 eV X-rays from the binding energy of the M-shell of the 68Ge/71Ge decays. Using the ratio of the K and L-shells, a total of 100 events per detector was expected after irradiation with 252Cf. Although an indication of such events is seen, no conclusive signal was yet observed due to the proximity to the noise edge and the lack of statistics.
@@
The linearity of the DAQ chain in the sub-keVee region was investigated using the pulse generator signals. The results are shown in Extended Data Fig. 5. Deviations from a pure linear behavior are observed below 250 eVee, where a few eVee variations have a strong impact on the CENS signal. They were attributed to two DAQ related effects [34]. The energy non-linearity was corrected in the CENS analysis and was taken into account during the event energy reconstruction.
@@
The energy resolution of the detectors at low energies was also evaluated with the pulse generator signals. For the C2 detector, a resolution of () eVee (FWHM) was found, while the C5 and C3 detectors have a resolution of () eVee (FWHM). An additional contribution to the total energy resolution comes from the statistical fluctuation of the number of electron-hole pairs produced in the Germanium crystals in the case of an event, which is
@@
with = 2.96 eV, the average energy needed to create a single electron-hole pair in Germanium, and the fano factor .
@@
Selection cuts are applied to reduce background events while keeping the CENS signal. Four different selection cuts are applied to the data: First, the muon-veto system allows for efficient suppression of the impact of cosmic radiation, using a 450 s anti-coincidence window between the veto and HPGe signals. The average rate detected in the muon veto during reactor ON is () Hz and decreases to () Hz in reactor off periods. The corresponding average dead times in both periods are 12.3% (reactor on) and 9.6% (reactor off). Second, inhibit signals are generated when the increasing baseline has reached saturation of the dynamic range of the transistor reset preamplifier (TRP) for each HPGe detector [28]. An anti-coincidence window of 1 to 2.5 ms (depending on the detector) is applied to veto unwanted spurious HPGe detector signals generated shortly after the resets. This cut suppressed 30% events after the muon veto anti-coincidence at low energy, becoming negligible above 5 keVee. The dead time induced by this cut is calculated combined with the previously mentioned muon-veto dead time to avoid the overlapping of both veto windows. An additional dead time of 0.5-2.1% (depending on the detector) is estimated. Third, the time difference distribution (TDD) of events is studied in each channel as proposed in [26]. Finally, an anti-coincidence cut is applied between different HPGe detectors with a 5 ms time window. The probability of a neutrino interacting with different detectors is negligible, while for other backgrounds, such as muon-induced neutrons created in the shield, simultaneous hits in several detectors at once can be expected.
The rejection efficiencies of these selection cuts are summarized in Extended Data Table 2 for the three detectors in different energy regions. The total dead time induced by the selection cuts is between 12.8% and 14.4% in the reactor on periods.
@@
The background model used in the analysis of the Conus+ data is based on Monte Carlo simulations using the Geant4-based framework MaGe [39] following the approach taken in [29]. A complete decomposition of the background in both reactor on and off data was achieved for all detectors used in the analysis. In the following, the main sources of background are described.
@@
As described in [30], the impact of cosmogenic neutrons with energies up to 100 MeV was investigated by first propagating the expected neutron spectrum at the KKL location (based on [40],[41]) with a flux of neutrons s-1cm-2 through a reactor building model. Then the resulting flux is tracked inside the Conus+ room. This flux was used as the basis for a next simulation, in which neutrons were started isotropically from a half-sphere around the Conus+ shield and their contribution to the Conus+ background was measured. The simulations show a contribution of counts d-1kg-1 in each detector in the energy range between 0.4 keVee and 1 keVee, which corresponds to approximately half of the background counts in this region.